# Matrix Multiplication in the MPC Model

### v1

@@ 3 Rectangular Matrix Multiplication
Observe that there would be a total of $n$ submatrices of the forms $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $n$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(d)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{ij}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$ be the set of processor assigned to $C^{ij}$. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\sqrt{d}+k$ i.e. in round 1, it receives $A^{1((i-1)\sqrt{d}+1)}$ and $B^{((i-1)\sqrt{d}+1)1}$, in round 2, it receives $A^{1((i-1)\sqrt{d}+2)}$ and $B^{((i-1)\sqrt{d}+2)1}$ and so on. Processor $p_{i}$ then computes the product of these submatrices locally after each round, adding the newly computed product to the existing product that it computed in the previous round using simple matrix addition. It’s easy to observe that after $\sqrt{d}$ rounds, processor $p_{i}$ will contain $\sum_{k=1}^{\sqrt{d}}A^{1((i-1)\sqrt{d}+k)}*B^{((i-1)\sqrt{d}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{ij}$, each containing a matrix of size $\sqrt{d}\times\sqrt{d}$ such that the sum of those matrices will produce $C^{11}$. According to Lemma 3.2, $C^{11}$ can be computed in $O(\log_{d}n)$ rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\sqrt{d}+\log_{d}n)$ rounds. The only thing missing in the proof is to show that each processor is sending messages of $O(d)$ size per round. Notice that each element of matrices $A$ and $B$ needs to be sent to at most $\sqrt{d}$ processors. Each processor initially holds $d$ elements of both $A$ and $B$. Thus, it needs to send $d\sqrt{d}$ elements to another processor, which can easily be done in $O(\sqrt{d})$ rounds, such that each processor sends $O(d)$ elements in each round. This finishes the proof for the upper bound.
Lower Bound:
The lower bound follows from Lemma 2.3 and Lemma 3.4. There are $d^{2}n$ terms that need to be calculated. In one round, a total of $O(nd\sqrt{d})$ terms can be calculated. Therefore, $\Omega(\sqrt{d})$ rounds will be required. We know that $c_{ij}=\Sigma_{k=1}^{n}a_{ik}b_{kj}$. Without loss of generality assume that all $b_{kj}=1$. Therefore, the processor computing $c_{ij}$ needs to compute the sum of $n$ elements that are initially present in at least $\frac{n}{d}$ processors because of $O(d)$ memory restriction. Hence, it will require $\Omega(\log_{d}n)$ rounds. Therefore, the final lower bound is $\Omega(\sqrt{d}+\log_{d}n)$
Theorem 3.6.
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $n$ processors of $O(d)$ memory requires $\Theta(\sqrt{d}+\log_{d}n)$ rounds in semirings, where $d\leq n$.
3.2.2 d processors with O(n) memory
@@
Observe that there would be a total of $d$ submatrices of the form $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $d$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(n)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{ij}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\frac{d}{\sqrt{n}}+k$ i.e. in round 1, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+1)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+1)1}$, in round 2, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+2)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+2)1}$ and so on. It is easy to observe that $p_{i}$ would receive these sub-matrices from a single other processor because of the way in which the initial distribution of $A^{iq}$ and $B^{qj}$ was done among the processors. Thus, the amount of information that a processor ends up sending in any communication round would be $O(n)$. Processor $p_{i}$ then computes the product of these submatrices locally after each round, adding the newly computed product to the existing product that it computed in the previous round using simple matrix addition. It’s easy to observe that after $\frac{d}{\sqrt{n}}$ rounds, processor $p_{i}$ will contain $\sum_{k=1}^{\frac{d}{\sqrt{n}}}A^{1((i-1)\frac{d}{\sqrt{n}}+k)}*B^{((i-1)\frac{d}{\sqrt{n}}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{ij}$, each containing a matrix of size $\sqrt{n}\times\sqrt{n}$ such that the sum of those matrices will produce $C^{11}$. According to Lemma 3.2, $C^{11}$ can be computed in $O(\log_{n}\frac{n}{d})$ rounds. But since $\frac{n}{d}\leq n$, it can be computed in a constant number of rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\frac{d}{\sqrt{n}})$ rounds. The lower bound follows from Lemma 2.3. There are $d^{2}n$ terms that need to be calculated. In one round, a total of $O(dn\sqrt{n})$ terms can be calculated. Therefore, $\Omega(\frac{d}{\sqrt{n}})$ rounds will be required. We proved the following theorem.
Theorem 3.7.
@@
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $d$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings, where $d\leq n$.
@@ 4 Sparse Matrix Multiplication
In this section, we will discuss the case when $A$ and $B$ are $d$-sparse. Notice that in this case, the resulting matrix $C=A*B$ can be $d^{2}$-sparse. But in the same spirit as discussed in Gupta et al. [4], in order to keep the model more meaningful, we assume that we are interested in only at most $d$ elements of each row and column of $C$. Therefore, both the input and output can be stored on $O(nd)$ memory. We analyse the case when we have $O(n)$ processors, each with $O(d)$ memory. We assume that initially each processor holds one row and one column (i.e. $O(d)$ elements in total) of $A$ and $B$, and at the end each processor outputs $d$ elements of one of the rows of $C$. We assume that the position of these $d$ elements of $C$ that a processor needs to output is already known to the processor.
@@
Notice each element of $C$ is the sum of $d$ terms. Thus, there are $nd^{2}$ terms that need to be calculated. In one round, a processor can compute any $d$ terms of $C$. It can do this by obtaining the $d$ elements of matrices $A$ and $B$ which are present in the terms, and then performing $d$ multiplications to compute the $d$ terms. Thus, in one round of local computation, the $n$ processors can compute any $nd$ terms of $C$. Hence, to compute the $nd^{2}$ of $C$, we only require $O(d)$ communication rounds. Now, we give a better algorithm than this as follows.
We try to find a more efficient method to compute the terms of $C$. For that, we draw inspiration for Gupta et al. [4]. In particular, we use Lemma 4.2 of their paper. As we discussed earlier, in their paper, they translated the problem of matrix multiplication to triangle processing; here we try to restate their results according to the terminology we used in this paper as follows.
Lemma 4.1 ([4]).
@@
$|T^{\prime}|=O(d^{2-\epsilon_{2}}n)$
Let us try to explain the lemma in simple words. The lemma basically states that the multiplication of two $n\times n$, $d$-sparse matrices ($C=A*B$) that contain $O(d^{2-\epsilon_{1}}n)$ terms to be computed can be divided into two parts $C^{\prime}$ and $C^{\prime\prime}$ such that
  • $C^{\prime}=A_{1}\times B_{1}+A_{2}\times B_{2}+\ldots+A_{l}\times B_{l}$ ($l$ is same as mentioned in the lemma) where $C_{i}=A_{i}\times B_{i}$ such that both $A_{i}$ and $B_{i}$ consist of $(\frac{n}{d})$ submatrices $A_{i}^{j}$ and $B_{i}^{j}$, for $j\in[\frac{n}{d}]$, of size $d\times d$ such that $A_{i}\times B_{i}=\sum A_{i}^{j}*B_{i}^{j}$
@@
Thus, the total time required will be the time to compute $C^{\prime}$ and $C^{\prime\prime}$. The idea is to use square matrix multiplication as a black box algorithm to compute $C^{\prime}$ and use a brute force algorithm to compute $C^{\prime\prime}$. To compute $C^{\prime}$, we can run $l$ sequential iterations, each for one $A_{i}\times B_{i}$. In one iteration, we can assign $d$ processors for the computation of $A_{i}^{j}*B_{i}^{j}$ for each $j$. These $(\frac{n}{d})$ computations will happen in parallel in one iteration. In this paper, we improve the bound on the number of iterations ($l$) required to go from $T$ to $T^{\prime}$. Notice that, in the MPC model, the time required for one iteration is $O(\sqrt{d})$.
@@ 4.1 Tight Bound for number of Iterations
We prove a tight upper bound on the number of iterations required to reach from $T$ to $T^{\prime}$(mentioned in Lemma 4.1).
Lemma 4.2.
Let $0\leq\varepsilon_{1}<\varepsilon_{2}$. Suppose $|T|\;=\;O(d^{2-\varepsilon_{1}}n)\quad\text{and}\quad|T^{\prime}|\;=\;O(d^{2-\varepsilon_{2}}n)$. Then the total number of iterations needed to reduce the number of uncomputed terms from $T$ to $T^{\prime}$ is $l\;=\;O\!\bigl{(}d^{4\,\varepsilon_{2}}\bigr{)}$.
Proof 4.3.
Let’s observe what happens when we introduce an intermediate step in the reduction of terms from $T$ to $T^{\prime}$. Let $0\leq\epsilon_{1}<\epsilon_{3}<\epsilon_{2}$ such that $\varepsilon_{3}\;=\;\frac{\varepsilon_{1}+\varepsilon_{2}}{2}$. Suppose we have
@@
Let $l_{1}$ denote the number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime\prime}$, and similarly, $l_{2}$ denote the number of iterations required to reduce the number of uncomputed terms from $T^{\prime\prime}$ to $T^{\prime}$. By Lemma 4.1, we have $l_{1}=O(d^{5\epsilon_{3}-\epsilon_{1}})$ and $l_{2}=O(d^{5\epsilon_{2}-\epsilon_{3}})$. Thus, the total number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime}$ is $l=l_{1}+l_{2}=O(d^{5\epsilon_{3}-\epsilon_{1}})+O(d^{5\epsilon_{2}-\epsilon_{3}})=O(d^{5\epsilon_{2}-\epsilon_{3}})$, (since $5\epsilon_{2}-\epsilon_{3}>{5\epsilon_{3}-\epsilon_{1}}$).
Now, if we recursively insert $\epsilon_{4}$ between $\epsilon_{3}$ and $\epsilon_{2}$ (such that $\epsilon_{4}$ is the arithmetic mean of $\epsilon_{3}$ and $\epsilon_{2}$) and similarly $\epsilon_{5}$ between $\epsilon_{4}$ and $\epsilon_{2}$ and do this $x$ number of times ($x$ is finite), then, we can see that number of iterations required to reduce the number of terms step by step would be $l=l_{1}+l_{2}+..+l_{x+1}$, where $l_{1}$ is the number of iterations to go from $\epsilon_{1}$ to $\epsilon_{3}$, $l_{2}$ from $\epsilon_{3}$ to $\epsilon_{4}$, .., $l_{x+1}$ from $\epsilon_{x+2}$ to $\epsilon_{2}$. Observe that the order of $l$ is dominated by the term $l_{x+1}$. Thus, $l=O(d^{5\epsilon_{2}-\epsilon_{x+2}})$. Hence, as $x$ becomes very large, $l$ converges to
$O(d^{5\epsilon_{2}-\epsilon_{x+2}})\approx O(d^{4\epsilon_{2}})$
@@
As we mentioned, the multiplication of two $d$-sparse matrices can be done in two phases. In one phase, we iteratively use square matrix multiplication, and in the other phase, we use the trivial algorithm. Our task is to compute $O(d^{2}n)$ terms. We will first reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\epsilon})$ ($\epsilon\geq 0$) by iteratively using square matrix multiplication, and then compute the rest of the terms using trivial computation. By Lemma 4.2, the number of iterations to reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\epsilon})$ is $O(d^{4\epsilon})$ where each iteration takes $O(\sqrt{d})$ communication rounds (complexity of square matrix multiplication). And computing $O(nd^{2-\epsilon})$ terms using trivial computation takes $O(\frac{nd^{2-\epsilon}}{nd})=O(d^{1-\epsilon})$communication rounds.Thus, the total number of communication rounds needed is $O(d^{4\epsilon+1/2})+O(d^{1-\epsilon})$, which is minimized at $\epsilon=0.1$ taking $O(d^{0.9})$ communication rounds. Thus, we prove the following theorem
Theorem 4.4.
Multiplication of two $d$-sparse matrices in the MPC model with $n$ processors, each with $O(d)$ memory, can be performed in $O(d^{0.9})$ rounds in semirings.

### v2

@@ 2 Rectangular Matrix Multiplication
Observe that there would be a total of $n$ submatrices of the forms $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $n$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(d)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{11}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$ be the set of processor assigned to $C^{11}$. We set the communication pattern as follows. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\sqrt{d}+k$. In other words, in round $1$, it receives matrices matrices $A^{1((i-1)\sqrt{d}+1)}$ and $B^{((i-1)\sqrt{d}+1)1}$, in round 2, it receives $A^{1((i-1)\sqrt{d}+2)}$ and $B^{((i-1)\sqrt{d}+2)1}$ and so on. In each round, $p_{i}$ computes the product of summatrices received in the current round, and adds the newly computed product to the existing matrix it computed till the previous round using simple matrix addition. It is easy to observe that after $\sqrt{d}$ rounds, processor $p_{i}$ will contain matrix $\sum_{k=1}^{\sqrt{d}}A^{1((i-1)\sqrt{d}+k)}*B^{((i-1)\sqrt{d}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{11}$, each containing a matrix of size $\sqrt{d}\times\sqrt{d}$ such that the sum of those matrices will produce $C^{11}$. According to lemma˜1, $C^{11}$ can be computed in $O(\log_{d}n)$ rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\sqrt{d}+\log_{d}n)$ rounds. The only thing missing in the proof is to show that each processor is sending messages of $O(d)$ size per round. Notice that each element of matrices $A$ and $B$ needs to be sent to at most $\sqrt{d}$ processors. Each processor initially holds $d$ elements of both $A$ and $B$. Thus, it needs to send $d\sqrt{d}$ elements to other processors, which can easily be done in $O(\sqrt{d})$ rounds, such that each processor sends $O(d)$ elements in each round. From this we get the following theorem.
Theorem 2.1
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $n$ processors of $O(d)$ memory can be done $O(\sqrt{d}+\log_{d}n)$ rounds in the MPC model, where $d\leq n$.
2.1.2 d processors with O(n) memory:
@@
Observe that there would be a total of $d$ submatrices of the form $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $d$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(n)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{11}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\frac{d}{\sqrt{n}}+k$ i.e. in round 1, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+1)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+1)1}$, in round 2, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+2)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+2)1}$ and so on. It is easy to observe that $p_{i}$ would receive these sub-matrices from a single other processor because of the way in which the initial distribution of $A^{iq}$ and $B^{qj}$ was done among the processors. Thus, the amount of information that a processor ends up sending in any communication round would be $O(n)$. Processor $p_{i}$ then computes the product of these submatrices locally after each round, adding the newly computed product to the existing product that it computed in the previous round using simple matrix addition. It’s easy to observe that after $\frac{d}{\sqrt{n}}$ rounds, processor $p_{i}$ will contain $\sum_{k=1}^{\frac{d}{\sqrt{n}}}A^{1((i-1)\frac{d}{\sqrt{n}}+k)}*B^{((i-1)\frac{d}{\sqrt{n}}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{ij}$, each containing a matrix of size $\sqrt{n}\times\sqrt{n}$ such that the sum of those matrices will produce $C^{11}$. According to lemma˜1, $C^{11}$ can be computed in $O(\log_{n}\frac{n}{d})$ rounds. But since $\frac{n}{d}\leq n$, it can be computed in a constant number of rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\frac{d}{\sqrt{n}})$ rounds.
Theorem 2.2
@@
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $d$ processors of $O(n)$ memory can be done in $O(\frac{d}{\sqrt{n}})$ rounds ounds in the MPC model, where $d\leq n$.
2.2 (n,d,n) Matrix Multiplicaton
In this case, we assume that we have $n$ processors with $O(n)$ memory each. Notice that in this case the matrices $A$ and $B$ can be stored in $O(nd)$ memory; however, the resulting $C$ matrix is a square matrix of size $n\times n$. Thus, we limit the total memory of the system to $O(n^{2})$. We divide the matrix $C$ into $n$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$ each, denoted by $C^{ij}$ (similar to section˜2.1) where $i,j\in[\sqrt{n}]$. Assign the computation of $C^{ij}$ to processor $p^{ij}$. Without loss of generality, assume that we scale the value of $d$ such that $d:=(\lceil\frac{d}{\sqrt{n}}\rceil\sqrt{n}).$ Divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{n}\times\sqrt{n}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,\ j\in[\sqrt{n}]$ and $q\in[\frac{d}{\sqrt{n}}]$. Observe that
We can assume that each processor $p_{ij}$ contains one row and one column of $A$ and $B$ respectively . Note that an element of $A$ needs to be sent to at most $\sqrt{n}$ processors. Thus, $p_{ij}$ needs to send $d\sqrt{n}$ elements. In one round, it can send $n$ elements and thus, in $\frac{d}{\sqrt{n}}$ rounds, it can send all the elements to the destination processors. Now let us analyse how many elements of $A$ and $B$ are received by each processor. $A^{iq}$ and $B^{qj}$ are $\sqrt{n}\times\sqrt{n}$ matrices containing a total of $n$ elements. Therefore, $p_{ij}$ needs a total of $(\frac{d}{\sqrt{n}}\cdot n)=d\sqrt{n}$ elements of $A$ and $B$ in order to compute $C^{ij}$. In one round, $p_{ij}$ can receive at most $O(n)$ elements and thus, in $\frac{d\sqrt{n}}{n}=\frac{d}{\sqrt{n}}$ rounds, it can receive all the elements that it requires for the computation of $C^{ij}$.
Theorem 2.3
Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ respectively, with $n$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds, where $d\leq n$.
@@ 3 Sparse Matrix Multiplication
In this section, we will discuss the case when $A$ and $B$ are $d$-sparse. Notice that in this case, the resulting matrix C=ABC=A*B can be $d^{2}$-sparse. But in the same spirit as discussed in Gupta et al. [7], in order to keep the model more meaningful, we assume that we are interested in only at most $d$ elements of each row and column of $C$. Therefore, both the input and output can be stored on $O(nd)$ memory. We analyse the case when we have $n$ processors, each with $O(d)$ memory. We assume that initially each processor holds one row and one column (i.e. $O(d)$ elements in total) of $A$ and $B$, and at the end each processor outputs $d$ elements of one of the rows of $C$. We assume that the position of these $d$ elements of $C$ that a processor needs to output is already known to the processor.
@@
Notice each element of $C$ is the sum of $d$ pairwise products of elements of $A$ and $B$. Thus, there are at most $nd^{2}$ products that need to be calculated. In one round, a processor can compute any $d$ products of $C$ that are required to compute an element of $C$. It can do this by obtaining the $d$ elements of matrices $A$ and $B$, then performing $d$ multiplications locally. Thus, in one round of local computation, the $n$ processors can compute any $nd$ terms of $C$. Hence, to compute the $nd^{2}$ of $C$, we only require $O(d)$ communication rounds. Now, we give a better algorithm than this as follows.
We try to find a more efficient method to compute the terms of $C$. For that, we adapt the idea of Gupta et al. [7]. In particular, we use Lemma 4.2 of their paper. As we discussed earlier, in their paper, they translated the problem of matrix multiplication to triangle processing. So first, we translate their results according to the terminology we used in this. We call each product of type $a_{ij}b_{jk}$ a term, where $a_{ij}$ and $b_{jk}$ are elements of $A$ and $B$ respectively. As we mentioned, there are at most $nd^{2}$ terms that needs to be computed inorder to compute $C$, Gupta et al. [7] computed them in a sequence (or layer) of rounds as follows.
Lemma 2([7])
@@
$|T^{\prime}|=O(d^{2-\varepsilon_{2}}n)$
Let us try to explain the lemma in simple words. The lemma basically states that the multiplication of two $n\times n$, $d$-sparse matrices (C=ABC=A*B) that contain $O(d^{2-\epsilon_{1}}n)$ terms to be computed can be divided into two parts $C^{\prime}$ and $C^{\prime\prime}$ such that
  1. 1.
    $C^{\prime}=A_{1}\times B_{1}+A_{2}\times B_{2}+\ldots+A_{l}\times B_{l}$ ($l$ is same as mentioned in the lemma) where
    1. (i)
      $A_{i}$ and $B_{i}$ are submatrices of $A$ and $B$ respectively.
    2. (ii)
      $A_{i}$ and $B_{i}$ consist of $(\frac{n}{d})$ submatrices $A_{i}^{j}$ and $B_{i}^{j}$, for $j\in[\frac{n}{d}]$, respectively. Each $A_{i}^{j}$ and $B_{i}^{j}$ is of size $d\times d$ such that $A_{i}\times B_{i}=\sum A_{i}^{j}*B_{i}^{j}$.
  2. 2.
@@
  3. 3.
    $C=C^{\prime}+C^{\prime\prime}$
Thus, the total time required to compute $C$ is the time to compute $C^{\prime}$ and $C^{\prime\prime}$ one after another. The idea is to use square matrix multiplication as a black box algorithm to compute $C^{\prime}$ (because computing $A_{i}$ and $B_{i}$ requires multiplication of many $d\times d$ square matrices) and use a brute force algorithm to compute $C^{\prime\prime}$.
To compute $C^{\prime}$, we can run $l$ sequential iterations, each for one $A_{i}\times B_{i}$. In one iteration, we can assign $d$ processors for the computation of $A_{i}^{j}*B_{i}^{j}$ for each $j$. These $(\frac{n}{d})$ computations will happen in parallel in one iteration. In this paper, we improve the bound on the number of iterations ($l$) required to go from $T$ to $T^{\prime}$. Notice that, in the MPC model, the time required for one iteration is $O(\sqrt{d})$. Because this involves $\frac{n}{d}$ instance matrix multiplications of $d\times d$ matrices happening in parallel, each with the help of $d$ computers. Thus, if we plug $n=d$ in theorem˜2.3, we get an $O(\sqrt{d})$ rounds algorithm for this task. Now we improve the bounds on the number of iterations required to bring down the number of terms from $O(d^{2-\epsilon_{1}})$ to $O(d^{2-\epsilon_{2}})$ mentioned in lemma˜2
@@ 3.1 Tight Bound for number of Iterations
Lemma 3
Number of rounds required to bring down the number of uncomputed terms from $|T|\;=\;O(d^{2-\varepsilon_{1}}n)$ to $|T^{\prime}|\;=\;O(d^{2-\varepsilon_{2}}n)$ is $l\;=\;O\!\bigl(d^{4\,\varepsilon_{2}}\bigr)$, where $0\leq\varepsilon_{1}<\varepsilon_{2}$.
Proof
According to lemma˜2 it will require $O(d^{5\varepsilon_{2}-\varepsilon_{1}})$ iteration to go from $T$ to $T^{\prime}$. Let us see what happens when we introduce an intermediate step in the reduction of terms from $T$ to $T^{\prime}$, i.e. instead of directly going from $O(d^{2-\varepsilon_{1}}n)$ to $O(d^{2-\varepsilon_{2}}n)$. We first go from $O(d^{2-\varepsilon_{1}}n)$ to $O(d^{2-\varepsilon_{3}}n)$ and then go from $O(d^{2-\varepsilon_{3}}n)$ to $O(d^{2-\varepsilon_{2}}n)$ for some $\varepsilon_{1}<\varepsilon_{3}<\varepsilon_{2}$, using lemma˜2. And calculate the number of iterations consumed in this process. Let $\varepsilon_{3}\;=\;\frac{\varepsilon_{1}+\varepsilon_{2}}{2}$ such that
@@
Let $l_{1}$ denote the number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime\prime}$, and similarly, $l_{2}$ denote the number of iterations required to reduce the number of uncomputed terms from $T^{\prime\prime}$ to $T^{\prime}$. By lemma˜2, we have $l_{1}=O(d^{5\varepsilon_{3}-\varepsilon_{1}})$ and $l_{2}=O(d^{5\varepsilon_{2}-\varepsilon_{3}})$. Thus, the total number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime}$ is $l=l_{1}+l_{2}=O(d^{5\varepsilon_{3}-\varepsilon_{1}})+O(d^{5\varepsilon_{2}-\varepsilon_{3}})=O(d^{5\varepsilon_{2}-\varepsilon_{3}})$. Since $5\varepsilon_{2}-\varepsilon_{3}>{5\varepsilon_{3}-\varepsilon_{1}}$, we get a strictly better bound than presented in lemma˜2.
Now the idea is that we recursively insert intermediate steps in this process i.e. insert $\varepsilon_{4}$ between $\varepsilon_{3}$ and $\varepsilon_{2}$ (such that $\varepsilon_{4}$ is the arithmetic mean of $\varepsilon_{3}$ and $\varepsilon_{2}$) and similarly $\varepsilon_{5}$ between $\varepsilon_{4}$ and $\varepsilon_{2}$ and do this $x$ number of times ( where $x$ is big constant), then, we can see that number of iterations required to reduce the number of terms step by step would be $l=l_{1}+l_{2}+..+l_{x+1}$, where $l_{1}$ is the number of iterations to go from $\varepsilon_{1}$ to $\varepsilon_{3}$, $l_{2}$ from $\varepsilon_{3}$ to $\varepsilon_{4}$, .., $l_{x+1}$ from $\varepsilon_{x+2}$ to $\varepsilon_{2}$. Observe that the order of $l$ is dominated by the term $l_{x+1}$. Thus, $l=O(d^{5\varepsilon_{2}-\varepsilon_{x+2}})$. Therefore, by inserting enough intermediate steps (bounded by some constant) we can make $\varepsilon_{x+2}$, infinitesimally $\varepsilon_{x+2}$, infinitesimally close to $\varepsilon_{2}$. Hence $l$ will converge to
$O(d^{5\varepsilon_{2}-\varepsilon_{x+2}})\approx O(d^{4\varepsilon_{2}})$
@@
As we mentioned, the multiplication of two $d$-sparse matrices can be done in two phases. In one phase, we iteratively use square matrix multiplication, and in the other phase, we use the brute-force algorithm. Our task is to compute $O(d^{2}n)$ terms. We will first reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\varepsilon})$ ($\varepsilon\geq 0$) by iteratively using square matrix multiplication, and then compute the rest of the terms using trivial computation. By lemma˜3, the number of iterations to reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\varepsilon})$ is $O(d^{4\varepsilon})$, where each iteration takes $O(\sqrt{d})$ communication rounds. Thus the total time for phase one is $O(d^{4\varepsilon+\frac{1}{2}})$. After phase one, we are left with $O(nd^{2-\varepsilon})$ terms that we compute using a trivial brute force algorithm. We know that in one round, one processor can compute $O(d)$ terms (because it has $O(d)$ memory), therefore total rounds required to compute $O(nd^{2-\varepsilon})$ temrs will be $O(\frac{nd^{2-\varepsilon}}{nd})=O(d^{1-\varepsilon})$. There is a tradeoff between the round complexity of these two phases, i.e. if we spend less time in phase one, it will take more time (polynomially) in phase two. Therefore, in order to optimise the total number of communication rounds needed, which is $O(d^{4\varepsilon+1/2})+O(d^{1-\varepsilon})$, should be minimised. Which happen for $\varepsilon=0.1$, taking $O(d^{0.9})$ communication rounds. Thus, we prove the following theorem.
Theorem 3.1
Multiplication of two $d$-sparse matrices in the MPC model with $n$ processors, each with $\Theta(d)$ memory, can be performed in $O(d^{0.9})$ rounds.